# Detecting LLM-Generated Korean Text through Linguistic Feature Analysis

### v2

@@ Abstract
The rapid advancement of large language models (LLMs) increases the difficulty of distinguishing between human-written and LLM-generated text. Detecting LLM-generated text is crucial for upholding academic integrity, preventing plagiarism, protecting copyrights, and ensuring ethical research practices. Most prior studies on detecting LLM-generated text focus primarily on English text. However, languages with distinct morphological and syntactic characteristics require specialized detection approaches. Their unique structures and usage patterns can hinder the direct application of methods primarily designed for English. Among such languages, we focus on Korean, which has relatively flexible spacing rules, a rich morphological system, and less frequent comma usage compared to English. We introduce KatFish, the first benchmark dataset for detecting LLM-generated Korean text. The dataset consists of text written by humans and generated by four LLMs across three genres.
By examining spacing patterns, part-of-speech diversity, and comma usage, we illuminate the linguistic differences between human-written and LLM-generated Korean text. Building on these observations, we propose KatFishNet, a detection method specifically designed for the Korean language. KatFishNet achieves an average of 19.78% higher AUROC compared to the best-performing existing detection method. Our code and data are available at https://github.com/Shinwoo-Park/detecting_llm_generated_korean_text_through_linguistic_analysis.
@@
Detecting LLM-Generated Korean Text through Linguistic Feature Analysis
footnotetext: $\dagger$ Corresponding author.
@@ 1 Introduction
2 Dataset Construction: KatFish
@@
We select these three genres to build a dataset that captures a diverse range of real-life scenarios and linguistic features while addressing practical challenges in detecting LLM-generated Korean text. Each genre highlights the importance of accurate detection in real-world applications.
@@ 2.2 LLM-Generated Text Collection
3 Detection Method: KatFishNet
@@ 3.1 Word Spacing Patterns
Unlike English, Korean has many exceptions and flexibilities in its spacing rules, making it one of the most variable and challenging aspects of writing. This makes it a valuable feature for examining stylistic and grammatical differences. In line with the Korean orthography word spacing guidelines [^4] https://korean.go.kr/kornorms/regltn/regltnView.do?regltn_code=0001&regltn_no=182#a182, we examine bound nouns (BN) and auxiliary predicate elements (VX), which are the key part-of-speech (POS) categories related to word spacing. We also investigate Eojeol POS Diversity and Unspaced VX Diversity in Appendix D. We use Bareun POS tagger.
Bound nouns must be used in conjunction with another word but require word spacing as they function as nouns. Their dependent nature leads to frequent mistakes, accounting for 70% of total word spacing errors (Shin et al., 2015). We examine two metrics: MMN-BN Space Ratio and BN Space Ratio. MMN-BN Space Ratio measures the frequency of word spacing between a numeral determiner (MMN) and a BN. BN Space Ratio quantifies the frequency of word spacing before a BN. To focus on cases with greater variation, we eliminate trivial cases where spaces may be omitted.
@@ 3.2 Part-of-Speech N-gram Diversity
Figure 3 compares the average POS n-gram diversity scores between human-written and LLM-generated essays. We include the analysis results for poetry and paper abstracts in the Appendix E due to space limitations. Human-written essays exhibit a higher diversity score than LLM-generated essays. The analysis results highlight that humans use a wider range of grammatical structures and construct sentences more flexibly than LLMs. Since LLMs generate text by selecting the most probable word combinations based on training data, they tend to repeat commonly used structures more frequently. {mdframed} Finding. Humans tend to use a more diverse range of POS combinations in their writing compared to LLMs.
@@ 3.3 Comma Usage Patterns
In Korean, commas help improve readability and clarify meaning within sentences. We analyze differences in comma usage to compare how humans and LLMs structure sentences and manage their flow. Specifically, we investigate the proportion of commas within sentences, their placement, structural changes in sentences, and linguistic diversity around commas. We compute the following five metrics: 1) Comma Inclusion Rate: The proportion of sentences containing at least one comma out of all sentences in a text. 2) Average Comma Usage Rate: The number of commas in a sentence divided by the total number of morphemes in that sentence. 3) Average Relative Position of Comma: The position of each comma (counting the number of morphemes before it) divided by the total number of morphemes in the sentence. If multiple commas appear in a sentence, the average relative position is calculated. 4) Average Segment Length: The average length of sentence segments split by commas. 5) POS Diversity Score Before and After Comma: The diversity of part-of-speech pairs appearing before and after a comma. This score is calculated by dividing the number of unique POS pairs by the total number of POS pairs. We segment each text into individual sentences, compute these metrics for each sentence, and use the average values of these sentence-level metrics as the representative values for the entire text. Appendix F further explores the specific POS patterns surrounding commas.
@@ 3.4 Design of KatFishNet
Our detection method, KatFishNet, is a machine learning-based model that leverages Korean linguistic features. KatFishNet offers several advantages: First, it enables analysis of which features play a crucial role in detecting LLM-generated text, making the model highly interpretable. Second, compared to deep learning-based models such as Transformer-based approaches, it does not require training text embeddings or large-scale datasets, allowing for a lightweight and efficient detection system. Lastly, traditional machine learning models like logistic regression, random forests, and support vector machines provide a practical and efficient solution by enabling training on CPUs, avoiding the costs associated with GPU resources.
@@ 4.1 Task Definition
We address a binary classification task where, given a Korean text, the goal is to classify whether the text is generated by a human or an LLM. We evaluate performance using the Area Under the Receiver Operating Characteristic curve (AUROC). AUROC measures the area under the curve that plots the True Positive Rate against the False Positive Rate at different threshold levels. This metric measures overall detection performance across all potential classification thresholds.
@@ 4.2 Baselines
The implementation details of KatFishNet and the baselines are provided in the Appendix G.
@@ 4.3 OOD Evaluation: Unseen LLMs
Table 3: Performance of detecting LLM-generated Korean text. We report the average performance (AUROC) over five experiments. We separately report the performance of the detection model for the task of distinguishing between human-written text and text generated by a specific LLM. For example, $\rightarrow$ Solar indicates that the detection model is evaluated on the task of classifying human-written text and text generated by Solar.
@@ 5 Experimental Results
We use logistic regression as the backbone model for KatFishNet and provide additional experimental results with random forest and support vector machine models in Table 8. The results show that KatFishNet achieves the highest performance when leveraging comma usage patterns, compared to spacing patterns and POS n-gram diversity. KatFishNet with comma usage patterns outperforms all other methods across all three text genres. It achieves a 16.74% performance improvement over LLM paraphrasing for essays, a 10.72% improvement over DetectGPT for poetry, and a 31.90% improvement over LLM paraphrasing for paper abstracts. Meanwhile, KatFishNet with POS n-gram diversity as features ranks second among all methods for essay, while KatFishNet with spacing patterns as features ranks second for abstract. We hypothesize that comma usage patterns are more useful for detecting LLM-generated Korean text than spacing patterns or POS n-gram diversity, as LLMs tend to have more difficulty learning comma usage than word spacing or POS combinations. Spacing follows relatively clear patterns in training data, and POS sequences can be learned as probabilistic patterns. In contrast, comma usage reflects contextual and stylistic factors, making it highly variable depending on the intent of the writer.
@@ 6 Related Work
Zellers et al. (2019) developed the GROVER dataset, which includes human-written and AI-generated news articles to support research on detecting machine-generated disinformation. Similarly, Fagni et al. (2021) created TweepFake, a dataset of tweets authored by both humans and bots, facilitating studies on social media content authenticity. Guo et al. (2023) introduced the HC3 dataset, which contains questions and answers generated by both human experts and ChatGPT in English and Chinese. Further advancing this line of research, Wang et al. (2024b) introduced the M4 dataset, a multi-generator, multi-domain, and multilingual corpus.
@@ 7 Conclusion
Appendix A Essay Topics for Argumentative Writing Across Education Levels
@@
Appendix B Prompt for Korean Text Generation via LLMs
@@
Essay: 너는 과제로 주장글을 작성해야 하는 [EDUCATION LEVEL] 학생이야.
@@
Essay: You are a [EDUCATION LEVEL] student who has to write an argumentative essay as an assignment.
@@
Poetry: You are a poet in your [AGE GROUP]. Read the given poem and understand its content.
@@
Paper Abstract: I’ll give you a paper without the abstract. Write an abstract for this paper.
@@
Appendix C Data Statistics: Eojeol Counts
@@
Appendix D Word Spacing Analysis
@@
Appendix E Part-of-Speech N-gram Diversity
Figure 6 compares the average POS n-gram diversity scores between human-written and LLM-generated Korean text. Excluding POS 4-gram and 5-gram in poetry, human-written text shows a higher diversity score than LLM-generated text. Poetry has shorter length and simpler structure compared to essays and paper abstracts, which can reduce the difference in POS n-gram diversity between human-written and LLM-generated textx for 4-gram and 5-gram sequences.
@@ Appendix E Part-of-Speech N-gram Diversity
Appendix F POS-Comma Analysis
@@ F.1 Before Comma
F.2 Before and After Comma
@@
Appendix G Implementation Details
@@
Appendix H Detection Results From the LLMs Used to Generate the KatFish Dataset
Table 7 presents the performance of the LLM-generated Korean text detection task using the LLMs employed in the creation of the KatFish dataset. The results suggest that even LLMs struggle to identify their own generated text, highlighting the limitations of LLM prompting baseline.
Essay GPT-4o 57.75 49.70 46.26 51.23
Solar 46.50 48.11 47.38 47.33
Qwen2 62.04 61.03 54.79 59.28
Llama3.1 52.64 37.95 45.11 45.23
Poetry GPT-4o 51.70 52.28 50.06 51.34
Solar 49.99 51.15 44.34 48.49
Qwen2 47.17 49.91 45.79 47.62
Llama3.1 64.93 48.58 50.60 54.70
Paper Abstract GPT-4o 48.20 47.70 47.77 47.89
Solar 51.49 52.32 53.05 52.28
Qwen2 50.00 50.50 49.52 50.00
Llama3.1 49.80 49.32 49.68 49.60
Table 7: Performance Comparison of Different LLMs in Detecting LLM-Generated Korean Text. We report the average performance (AUROC) over five experiments. We separately report the performance of the detection model for the task of distinguishing between human-written text and text generated by a specific LLM. For example, $\rightarrow$ Solar indicates that the detection model is evaluated on the task of classifying human-written text and text generated by Solar.
Appendix I Performance of KatFishNet with Different Machine Learning Models
Table 8 presents the performance of KatFishNet with different base machine learning models. We find that KatFishNet, which leverages comma usage patterns, achieves the highest performance in detecting LLM-generated Korean text, regardless of the type of machine learning model used as the backbone.
@@ Appendix I Performance of KatFishNet with Different Machine Learning Models
Table 8: Performance of detecting LLM-generated Korean text. We report the average performance (AUROC) over five experiments.
Appendix J Methods for Detection Machine-Generated Text
Methods for detecting machine-generated text can be broadly grouped into three main categories:
@@
Watermarking methods (Kirchenbauer et al., 2023; Guo et al., 2024) detect machine-generated text by embedding recognizable patterns during the text generation process. These methods intervene in the text generation process of LLMs by dividing the vocabulary into green and red lists and prioritizing the generation of tokens from the green list. If the proportion of green tokens in a generated text exceeds a certain threshold, the text is classified as machine-generated. Since these methods require direct manipulation of the decoding process to embed watermarks, their application is limited to open-source LLMs.

### v4

# Shinwoo Park  Shubin Kim  Do-Kyung Kim  Yo-Sub Han

@@ Abstract
The rapid advancement of large language models (LLMs) increases the difficulty of distinguishing between human-written and LLM-generated text. Detecting LLM-generated text is crucial for upholding academic integrity, preventing plagiarism, protecting copyrights, and ensuring ethical research practices. Most prior studies on detecting LLM-generated text focus primarily on English text. However, languages with distinct morphological and syntactic characteristics require specialized detection approaches. Their unique structures and usage patterns hinder the direct application of methods primarily designed for English. Among such languages, we focus on Korean, which has relatively flexible spacing rules, a rich morphological system, and less frequent comma usage compared to English. We introduce KatFish, the first benchmark dataset for detecting LLM-generated Korean text. The dataset consists of text written by humans and generated by four LLMs across three genres. By examining spacing patterns, part-of-speech diversity, and comma usage, we illuminate the linguistic differences between human-written and LLM-generated Korean text. Building on these observations, we propose KatFishNet, a detection method specifically designed for the Korean language. KatFishNet achieves an average of 19.78% higher AUC-ROC compared to the best-performing existing detection method. Our code and data are available at https://github.com/Shinwoo-Park/katfishnet.
@@
KatFishNet: Detecting LLM-Generated Korean Text through
Linguistic Feature Analysis
footnotetext: $\star$ Corresponding author.
@@ 1 Introduction
2 Dataset Construction: [Uncaptioned image]KatFish
@@
We select these three genres to build a dataset that captures a diverse range of real-life scenarios and linguistic features while addressing practical challenges in detecting LLM-generated Korean text. Each genre highlights the importance of accurate detection in real-world applications. For a detailed rationale behind our genre selection, see Appendix A.
@@ 2.2 LLM-Generated Text Collection
3 Detection Method: [Uncaptioned image]KatFishNet
@@ 3.1 Word Spacing Patterns
Unlike English, Korean has many exceptions and flexibilities in its spacing rules, making it one of the most variable and challenging aspects of writing. This makes it a valuable feature for examining stylistic and grammatical differences. In line with the Korean orthography word spacing guidelines [^4] https://korean.go.kr/kornorms/regltn/regltnView.do?regltn_code=0001&regltn_no=182#a182, we examine bound nouns (BN) and auxiliary predicate elements (VX), which are the key part-of-speech (POS) categories related to word spacing. We also investigate Eojeol POS Diversity and Unspaced VX Diversity in Appendix E. We use Bareun POS tagger.
Bound nouns must be used in conjunction with another word but require word spacing as they function as nouns. Their dependent nature leads to frequent mistakes, accounting for 70% of total word spacing errors (Shin et al., 2015). We examine two metrics: Numeral Determiner–Bound Noun (MMN-BN) Space Ratio and Bound Noun (BN) Space Ratio. The MMN-BN Space Ratio measures the frequency of word spacing between a numeral determiner (MMN) and a bound noun (BN). BN Space Ratio quantifies the frequency of word spacing before a BN. To focus on cases with greater variation, we eliminate trivial cases where spaces may be omitted.
@@ 3.2 Part-of-Speech N-gram Diversity
Figure 3 compares the average POS n-gram diversity scores between human-written and LLM-generated essays. We include the analysis results for poetry and paper abstracts in the Appendix F due to space limitations. Human-written essays exhibit a higher diversity score than LLM-generated essays. The analysis results highlight that humans use a wider range of grammatical structures and construct sentences more flexibly than LLMs. Since LLMs generate text by selecting the most probable word combinations based on training data, they tend to repeat commonly used structures more frequently. {mdframed} Finding. Humans tend to use a more diverse range of POS combinations in their writing compared to LLMs.
@@ 3.3 Comma Usage Patterns
In Korean, commas help improve readability and clarify meaning within sentences. We analyze differences in comma usage to compare how humans and LLMs structure sentences and manage their flow. Specifically, we investigate the proportion of commas within sentences, their placement, structural changes in sentences, and linguistic diversity around commas. We compute the following five metrics: 1) Comma Inclusion Rate: The proportion of sentences containing at least one comma out of all sentences in a text. 2) Average Comma Usage Rate: The number of commas in a sentence divided by the total number of morphemes in that sentence. 3) Average Relative Position of Comma: The position of each comma (counting the number of morphemes before it) divided by the total number of morphemes in the sentence. If multiple commas appear in a sentence, the average relative position is calculated. 4) Average Segment Length: The average length of sentence segments split by commas. 5) POS Diversity Score Before and After Comma: The diversity of part-of-speech pairs appearing before and after a comma. This score is calculated by dividing the number of unique POS pairs by the total number of POS pairs. We segment each text into individual sentences, compute these metrics for each sentence, and use the average values of these sentence-level metrics as the representative values for the entire text. Appendix G further explores the specific POS patterns surrounding commas.
@@ 3.4 Design of KatFishNet
Our detection method, KatFishNet, is a machine learning-based model that leverages Korean linguistic features. Table 8 in Appendix H presents the Korean linguistic features used by KatFishNet. KatFishNet offers several advantages: First, it enables analysis of which features play a crucial role in detecting LLM-generated text, making the model highly interpretable. Second, compared to deep learning-based models such as Transformer-based approaches, it does not require training text embeddings or large-scale datasets, allowing for a lightweight and efficient detection system. Lastly, traditional machine learning models like logistic regression, random forests, and support vector machines provide a practical and efficient solution by enabling training on CPUs, avoiding the costs associated with GPU resources.
@@ 4.1 Task Definition
We address a binary classification task where, given a Korean text, the goal is to classify whether the text is generated by a human or an LLM. We evaluate performance using the Area Under the Receiver Operating Characteristic curve (AUC-ROC). AUC-ROC measures the area under the curve that plots the True Positive Rate against the False Positive Rate at different threshold levels. This metric measures overall detection performance across all potential classification thresholds.
@@ 4.2 Baselines
The implementation details of KatFishNet and the baselines are provided in the Appendix I.
@@ 4.3 OOD Evaluation: Unseen LLMs
Table 3: Performance of detecting LLM-generated Korean text. We report the average performance (AUC-ROC) over five experiments. We separately report the performance of the detection model for the task of distinguishing between human-written text and text generated by a specific LLM. For example, $\rightarrow$ Solar indicates that the detection model is evaluated on the task of classifying human-written text and text generated by Solar.
@@ 5 Experimental Results
We use logistic regression as the backbone model for KatFishNet and provide additional experimental results with random forest and support vector machine models in Appendix J. The results show that KatFishNet achieves the highest performance when leveraging comma usage patterns, compared to spacing patterns and POS n-gram diversity. KatFishNet with comma usage patterns outperforms all other methods across all three text genres. It achieves a 16.74% performance improvement over LLM paraphrasing for essays, a 10.72% improvement over DetectGPT for poetry, and a 31.90% improvement over LLM paraphrasing for paper abstracts. Meanwhile, KatFishNet with POS n-gram diversity as features ranks second among all methods for essay, while KatFishNet with spacing patterns as features ranks second for abstract. We hypothesize that comma usage patterns are more useful for detecting LLM-generated Korean text than spacing patterns or POS n-gram diversity, as LLMs tend to have more difficulty learning comma usage than word spacing or POS combinations. Spacing follows relatively clear patterns in training data, and POS sequences can be learned as probabilistic patterns. In contrast, comma usage reflects contextual and stylistic factors, making it highly variable depending on the intent of the writer.
We analyze the detection performance of KatFishNet in Appendix K by combining three categories of linguistic features using an ensemble approach.
@@ 6 Related Work
Zellers et al. (2019) developed the GROVER dataset, which includes human-written and AI-generated news articles to support research on detecting machine-generated disinformation. Similarly, Fagni et al. (2021) created TweepFake, a dataset of tweets authored by both humans and bots, facilitating studies on social media content authenticity. Guo et al. (2023) introduced the HC3 dataset, which contains questions and answers generated by both human experts and ChatGPT in English and Chinese. Wang et al. (2024c) advanced this direction by introducing the M4 dataset, a multi-generator, multi-domain, and multilingual corpus.
Wang et al. (2024b) introduced M4GT-Bench, a multilingual, multi-domain corpus covering various LLMs and proposes more nuanced tasks, such as pinpointing the specific LLM behind the generated text (multi-way detection) and identifying human–machine boundary points in mixed-content texts. Similarly, Macko et al. (2023) developed MULTITuDE, a large-scale multilingual dataset spanning eleven languages, to examine the cross-lingual and cross-generator generalization capabilities of detectors, comparing multilingual detectors to monolingual detectors. By systematically varying the training and evaluation languages, MULTITuDE demonstrates that English-focused models often do not seamlessly transfer to other languages without explicit multilingual training.
Although M4GT-Bench and MULTITuDE both address multilingual machine-generated text detection, neither benchmark focuses on language-specific approaches. In M4GT-Bench, the monolingual setup is trained exclusively on English data, while the multilingual setup uses a mixture of languages excluding the target language. As a result, there is no training that captures the morphological or syntactic characteristics of any single language in particular. MULTITuDE trains a separate model for each language and applies the same detection method across all of them. Its primary goal is to explore language transfer—that is, to examine how a model trained on certain languages generalizes to others—rather than to enhance performance for individual languages.
In contrast, our work highlights Korean-specific features—such as flexible spacing rules, part-of-speech n-gram patterns, and punctuation usage—in the design of the detection method. By focusing on language-specific characteristics, we demonstrate how linguistically tailored strategies can enhance the performance of LLM-generated text detection.
@@ 7 Conclusion
Acknowledgments
This research was supported by the National Research Foundation of Korea (NRF) grants funded by the Korean government (MSIT) (RS-2025-02222626 and RS-2025-00562134), and by the AI Graduate School Program (RS-2020-II201361).
Appendix A Genre Selection Criteria and Rationale
Practicality-Driven Domain Selection: Targeting Directly Problematic Use Cases
Our selection of text types is driven by a focus on practical use cases where the use of LLMs for text generation is directly problematic. This contrasts with other genres where LLM use is either legitimate or not explicitly prohibited, but may still lead to indirect issues. We prioritize three genres that directly relate to violations of academic integrity, copyright, and research ethics—major issues that are of immediate concern in real-world applications of LLM-generated text.
Ensuring Broad Coverage in Writing Types
Texts are classified into different types by their purpose, writing style, and usage. To ensure the broad applicability of our findings, we meticulously choose three genres—student-written argumentative essays, free-verse poetry, and research paper abstracts—to represent major facets of writing. Table 4 shows how each genre in KatFish contributes to this coverage.
Category Type Genres
Purpose Expressive Writing Poetry
Informative Writing Paper Abstract
Persuasive Writing Essay
Writing Style Descriptive Poetry
Narrative Poetry, Essay
Expository Paper Abstract
Persuasive Essay
Usage Literary Writing Poetry
Academic Writing Essay, Paper Abstract
Table 4: Summary of genre characteristics in KatFish, illustrating their complementary roles across purpose, writing style, and usage.
We choose free-verse poetry, a less conventional representative for literary text, to investigate how less conventional writing affects LLM-generated text detection. This allows us to gain insights that cannot be obtained from more structured forms.
Appendix B Essay Topics for Argumentative Writing Across Education Levels
@@
Appendix C Prompt for Korean Text Generation via LLMs
@@
Essay 너는 과제로 주장글을 작성해야 하는 [EDUCATION LEVEL] 학생이야.
@@
Essay You are a [EDUCATION LEVEL] student who has to write an argumentative essay as an assignment.
@@
Poetry You are a poet in your [AGE GROUP]. Read the given poem and understand its content.
@@
Paper Abstract I’ll give you a paper without the abstract. Write an abstract for this paper.
@@
Appendix D Data Statistics: Eojeol Counts
@@
Appendix E Word Spacing Analysis
@@
Appendix F Part-of-Speech N-gram Diversity
Figure 6 compares the average POS n-gram diversity scores between human-written and LLM-generated Korean text. Excluding POS 4-gram and 5-gram in poetry, human-written text shows a higher diversity score than LLM-generated text. Poetry has shorter length and simpler structure compared to essays and paper abstracts, which can reduce the difference in POS n-gram diversity between human-written and LLM-generated texts for 4-gram and 5-gram sequences.
@@ Appendix F Part-of-Speech N-gram Diversity
Appendix G POS-Comma Analysis
@@ G.1 Before Comma
G.2 Before and After Comma
@@
Appendix H Summary of Linguistic Features
This section presents a summary of the linguistic features employed in KatFishNet. The feature set falls into three primary categories: word spacing, POS combinations, and punctuation. Table 8 outlines each feature with a brief description.
Category Feature Description
Word Spacing MMN-BN Space Ratio The frequency of word spacing between a MMN and a BN.
BN Space Ratio The frequency of word spacing before a BN.
VX Space Ratio The frequency of word spacing before a VX, excluding the specific case of "-아"/"-어" (ENDING) + "지" (VX), where spacing is strictly prohibited.
POS Combinations N-gram Diversity The syntactic variability in a text by computing the ratio of unique POS n-grams to the total number of POS n-grams.
Punctuation Comma Inclusion Rate The proportion of sentences containing at least one comma out of all sentences in a text.
Comma Usage Rate The number of commas in a sentence divided by the total number of morphemes in that sentence.
Relative Position The position of each comma (counting the number of morphemes before it) divided by the total number of morphemes in the sentence.
Segment Length The average length of sentence segments split by commas.
POS Diversity Score The diversity of part-of-speech pairs appearing before and after a comma.
Table 8: Overview of linguistic features used by KatFishNet across three categories.
Appendix I Implementation Details
@@
Appendix J Performance of KatFishNet with Different Machine Learning Models
Table 9 presents the performance of KatFishNet with different base machine learning models. Below is a brief overview of the models incorporated in KatFishNet:
  • Logistic Regression: A statistical model used when the dependent variable is categorical. It applies a sigmoid function to the input values to compute probabilities, and the regression coefficients help interpret how each feature influences the outcome.
  • Random Forest: An ensemble method that constructs multiple decision trees and combines their predictions to produce the final output. Each tree is trained on randomly selected subsets of the data and features, thus enhancing the diversity and generalization capability of the model.
  • Support Vector Machine: A classification algorithm that seeks an optimal decision boundary to separate data points. It learns by maximizing the margin between the decision boundary and the closest data points. This approach is effective in high-dimensional spaces and can handle complex data structures well.
We find that KatFishNet, which leverages comma usage patterns, achieves the highest performance in detecting LLM-generated Korean text, regardless of the type of machine learning model used as the backbone. Looking ahead, we also plan to develop an expanded version of KatFishNet based on ensemble or voting methods that integrate predictions from a diverse set of machine learning models. By understanding and leveraging the unique strengths of each model, we anticipate that our approach will achieve even stronger detection performance and robustness.
@@
Table 9: Performance of detecting LLM-generated Korean text. We report the average performance (AUC-ROC) over five experiments.
Appendix K Ensemble of Linguistic Features
We conduct an additional experiment to explore the effectiveness of combining word spacing, POS combinations, and punctuation features for LLM-generated Korean text detection. Specifically, we apply an ensemble approach by averaging the predicted probabilities (ranging from 0 to 1) from separate logistic regression models. A value of 1 indicates that the text is predicted to be LLM-generated. Tables 1011, and 12 show the experimental results.
Feature Combination $\rightarrow$ Solar $\rightarrow$ Qwen2 $\rightarrow$ Llama3.1 Average
Word Spacing 86.00 80.63 71.91 79.51
POS Combinations 92.26 83.10 73.63 82.99
Punctuation 97.57 94.63 92.45 94.88
All 97.67 93.93 88.05 93.21
POS + Punctuation 98.97 95.09 90.58 94.88
POS + Spacing 90.29 83.65 73.73 82.55
Punctuation + Spacing 97.89 94.74 90.35 94.32
Table 10: Performance on LLM-generated essay detection using individual and ensemble combinations of linguistic features. The ensemble method averages outputs from separate logistic regression classifiers.
Feature Combination $\rightarrow$ Solar $\rightarrow$ Qwen2 $\rightarrow$ Llama3.1 Average
Word Spacing 71.85 65.56 43.81 60.40
POS Combinations 39.41 79.17 53.32 57.30
Punctuation 62.65 93.45 63.22 73.10
All 63.66 96.04 59.94 73.21
POS + Punctuation 55.30 94.83 63.40 71.17
POS + Spacing 52.09 80.27 48.69 60.35
Punctuation + Spacing 73.78 95.41 58.64 75.94
Table 11: Performance on LLM-generated poetry detection using individual and ensemble combinations of linguistic features. The ensemble method averages outputs from separate logistic regression classifiers.
Feature Combination $\rightarrow$ Solar $\rightarrow$ Qwen2 $\rightarrow$ Llama3.1 Average
Word Spacing 57.73 66.91 49.36 58.00
POS Combinations 47.47 70.05 42.47 53.33
Punctuation 78.99 77.47 70.41 75.62
All 68.93 80.00 57.01 68.64
POS + Punctuation 71.70 82.82 60.93 71.81
POS + Spacing 53.22 70.76 44.98 56.32
Punctuation + Spacing 77.81 79.70 65.63 74.38
Table 12: Performance on LLM-generated paper abstract detection using individual and ensemble combinations of linguistic features. The ensemble method averages outputs from separate logistic regression classifiers.
In most cases, punctuation alone achieves detection performance that is either higher or comparable to that of ensemble models. This reinforces the strong effectiveness of punctuation-related features in detecting LLM-generated Korean text. Combining POS and spacing leads to the lowest performance among the ensembles, suggesting that these features alone do not capture the high-impact signals that punctuation features provide. Building on these findings, we plan to explore ensemble or voting approaches using multiple machine learning models, with a particular emphasis on punctuation features.
Appendix L Methods for Detection Machine-Generated Text
This section serves as supplementary material to Section 6, providing an overview of methodologies for detecting LLM-generated text. Methods for detecting machine-generated text can be broadly grouped into three main categories:
@@
Watermarking methods (Kirchenbauer et al., 2023; Guo et al., 2024) detect machine-generated text by embedding recognizable patterns during the text generation process. These methods intervene in the text generation process of LLMs by dividing the vocabulary into green and red lists and prioritizing the generation of tokens from the green list. If the proportion of green tokens in a generated text exceeds a certain threshold, the text is classified as machine-generated.